# Orthogonal projectors of binary LCD codes

### v1

@@ 1 Introduction
The interplay between codes and designs have provided many useful results [Assmus1992-rb]. Concerning relations between codes and graphs, several results are known: The rank of an adjacency matrix $A$ over $\mathbb{F}_{p}$ is called the $p$-rank of $A$. Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of strongly regular graphs are characterized by their eigenvalues. Haemers, Peeters and Rijckevorsel [Haemers1999-no] studied binary codes from adjacency matrices of strongly regular graphs. For strongly regular graphs with certain parameters, they observed that non-isomorphic graphs give inequivalent codes. Massey [massey_linear_1992] introduced linear complementary dual codes (LCD codes for short), and showed that a linear code has the orthogonal projector if and only if it is an LCD code. Since any orthogonal projector is symmetric matrix, adjacency matrices of some undirected graphs should coincide with orthogonal projectors. Based on this observation, Rodrigues and Keys studied a relation between LCD codes and adjacency matrices of graphs. They gave two conditions where adjacency matrices of graphs generate LCD codes over $\mathbb{F}_{q}$, where $q$ is a prime power [Key2018-ec, Proposition 2].
In this paper, we consider binary LCD codes in particular, and study a relation between binary LCD codes and adjacency matrices of simple graphs. It is known that the orthogonal projector $A$ of a binary LCD code is a symmetric matrix with $A^{2}=A$, where we regard $A$ as a matrix over $\mathbb{F}_{2}$. Hence any adjacency matrix satisfies the latter condition is an orthogonal projector of binary LCD code, as shown in [Key2018-ec, Brouwer1992-nc]. We show that such adjacency matrix is the orthogonal projector of a binary even LCD code, and two binary even LCD codes are equivalent if and only if the corresponding graphs are isomorphic. This establishes a one-to-one correspondence between binary even LCD codes and the above-mentioned class of graphs. Furthermore, we give a lower bound on any linear code obtained as a row span of strongly regular graphs. As an application, we consider codes generated by $srg(41,20,9,10)$, and partially solve a conjecture by Haemers, Peeters and Rijckevorsel [Haemers1999-no]. Furthermore, we improve some lower bounds on the minimum weight of binary LCD codes by constructing optimal binary $[n,k,d]$ LCD codes with the following parameters:
@@
We note that orthogonal projectors of the above optimal codes are adjacency matrices of regular graphs. By the above example, we show that from optimal binary LCD codes we can generate graphs whose adjacency matrices give codes with large minimum weight.
@@ 2 Preliminaries
Let $\mathbb{F}_{q}$ be the finite field of order $q$, where $q$ is a prime power. Let $\mathbb{F}_{q}^{n}$ be the vector space of all $n$-tuples over $\mathbb{F}_{q}$. A $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ is said to be an $[n,k]$ code over $\mathbb{F}_{q}$. Especially, codes over $\mathbb{F}_{2}$ are said to be binary codes. Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$. The parameters $n$ and $k$ are said to be the length and the dimension of $C$ respectively. A vector in $C$ is called a codeword. The weight of $x=(x_{1},x_{2},\dots,x_{n})\in\mathbb{F}_{q}^{n}$ is defined as $\operatorname{wt}(x)=|\{i\mid x_{i}\neq 0\}|$. The minimum weight $d(C)$ of $C$ is the minimum weight among all nonzero codewords in $C$. If the minimum weight of $C$ equals to $d$, then $C$ is said to be an $[n,k,d]$ code over $\mathbb{F}_{q}$. We say that a code $C$ is even if all codewords have even weights; otherwise, the code $C$ is odd. Two $[n,k]$ codes $C_{1}$ and $C_{2}$ over $\mathbb{F}_{q}$ are equivalent if there is a monomial matrix $M$ such that $C_{2}=\{cM\mid c\in C_{1}\}$. The equivalence of two codes $C_{1}$ and $C_{2}$ is denoted by $C_{1}\simeq C_{2}$. A generator matrix of a code $C$ is any matrix whose rows form a basis of $C$.
@@
Let $C$ be a code over $\mathbb{F}_{q}$ with length $n$ and let $T\subseteq\{1,2,\dots,n\}$. Deleting every coordinate $i\in T$ in every codeword of $C$ gives a code called the punctured code on $T$ and denoted by $C^{T}$. Define a subcode $C(T)=\{(c_{1},c_{2},\dots,c_{n})\in C\mid c_{i}=0\text{ for all }i\in T\}$. Puncturing $C(T)$ on $T$ gives a code called the shortened code on $T$ and denoted by $C_{T}$. If $|T|=1$, say $T=\{i\}$, then we will write $C^{\{i\}}$ and $C_{\{i\}}$ as $C^{i}$ and $C_{i}$, respectively. Bouyuklieva [bouyuklieva_optimal_2021, Proposition 1] showed that any binary LCD $[n,k,d]$ code can be extended to binary even LCD $[n+1,k,d+1]$ code provided $k$ is even. This result is based on the characterization by Carlet, Mesnager, Tang and Qi [carlet_new_2019, Theorem 3]. Furthermore, Bouyuklieva gave a characterization of both punctured codes and shortened codes of binary LCD codes:
@@
The rank of an adjacency matrix $A$ over $\mathbb{F}_{p}$ is called the $p$-rank of $A$. Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of strongly regular graphs are characterized by their eigenvalues. Haemers, Peeters and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs, and obtained many computational results for strongly regular graphs with small vertices. Keys and Rodrigues [Key2018-ec] constructed many LCD codes over $\mathbb{F}_{q}$ from adjacency matrices of strongly regular graphs via orthogonal projectors. Especially, they introduced the concept of reflexive LCD codes and complementary LCD codes [Key2018-ec, Definition 2], and provided the condition where adjacency matrices of graphs give LCD codes over $\mathbb{F}_{q}$ [Key2018-ec, Proposition 2].
@@ 3 Main results
In this section, we give main results on a correspondance between binary even LCD codes and adjacency matrices of certain simple graphs. Also, we provide a lower bound on the minimum weights of binary codes that come from strongly regular graphs.
@@ 3.1 Characterization
Let $e_{i}$ be the unit vector whose componets are all zero except $i$-th component, and let $r_{i}$ be the $i$-th row of $A$. Then, $Ae_{i}=r_{i}\in C$ for all $1\leq i\leq n$. Hence the row span of $A$ is a subspace of $C$. On the other hand, for any $c\in C$ it holds that $Ac=c$. This tells us that any codeword $c$ can be expressed as a linear combination of the rows of $A$. Hence $C$ is a subspace of the row span of $A$. This completes the proof. ∎
@@
  • If two vertices $x$ and $y$ are adjacent, then the number of the commmon neighbors of $x$ and $y$ is odd.
  • If two vertices $x$ and $y$ are not adjacent, then the number of the commmon neighbors of $x$ and $y$ is even.
@@
Bouyuklieva studied punctured codes and shortened codes of binary LCD codes and obtained Theorem 2.2. To prove the theorem, they essentially considered the orthogonal projection of the unit vector $e_{i}$ and made the following observation: If the $i$-th component of the orthogonal projection of $e_{i}$ is $1$, then $C_{i}$ is LCD and $C^{i}$ is not LCD; otherwise, $C^{i}$ is LCD and $C_{i}$ is not LCD. The $i$-th component of the orthogonal projection of $e_{i}$ is equal to the $i$-th diagonal entry of the orthogonal projector $A$. Hence it follows that $(A)_{i,i}=0$ if and only if $C^{i}$ is LCD, and $A_{i,i}=1$ if only if $C_{i}$ is LCD.
@@ 3.2 Bounds On Minimum Weights
In this section, we give a lower bound on the minimum weights of binary codes obtained as row spans of adjacency matrices of strongly regular graphs. These results will be applied in the following sections. Note that all the assertions in this section hold not only for binary LCD codes but also for any binary codes.
@@
However, this is absurd since the left hand side of the equation is even, while the right hand side is odd. This completes the proof. ∎
@@ 4 Applications
4.1 Binary Codes from $srg(41,20,9,10)$
Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of strongly regular graphs are characterized by their eigenvalues. For strongly regular graphs $srg(41,20,9,10)$, any corresponding binary code has $2$-rank, namely, dimension $20$. Haemers, Peeters and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs. They observed that adjacency matrices of known non-isomorphic $srg(41,20,9,10)$ gave inequivalent codes, and among such codes the code from the Paley graph had the largest minimum weight. They conjectured that the latter statement holds in general.
Remarkably, the former statement holds in general for any strongly regular graph with parameters $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ for $q\equiv 1\pmod{8}$ since any adjacency matrix of such a graph satisfies the condition given in Theorem 3.7. Furthermore, we can show that the Paley graph gives a binary even $[40,20]$ code by the following argument: The Griesmer bound shows that the largest minimum weight of binary LCD $[41,20]$ code is $11$. By Lemma 3.6, the minimum weight must be even. This proves that the Paley graph indeed gives a code with largest minimum weight among all $srg(41,20,9,10)$. Note that the classification of $srg(41,20,9,10)$ has not been completed yet. We state this fact as the following theorem:
Theorem 4.1.
Any pair of non-isomorphic $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q\equiv 1\pmod{8}$ gives a pair of inequivalent binary even LCD codes, and vice versa. Furthermore, the Paley graph of order $41$ gives a binary even $[40,20,10]$ code whose minimum weight is largest among all $srg(40,20,9,10)$.
@@ 4.1 Binary Codes from $srg(41,20,9,10)$
By Proposition 3.10, any code from $srg(41,20,9,10)$ must have the minimum weight larger than $d\geq 1+(k+1)/(\min\{\lambda,\mu\}+1)=3$. Applying Theorem 3.7, we can improve the lower bound by $1$, and hence the lower bound on the minimum weight is $4$. The above observation shows this lower bound is tight for $srg(40,20,9,10)$.
4.2 Construction by Orthogonal Projectors
In this section, we improve some lower bounds on the minimum weight of binary LCD codes by constructing optimal binary LCD codes from their orthogonal projectors. For recent results on optimal binary LCD codes, we refer the reader to [bouyuklieva_optimal_2021]. Let $d(n,k)$ denote the largest minimum weight among all $[n,k]$ LCD code.
@@ 4.2 Construction by Orthogonal Projectors
For each of these codes, the first rows of their orthogonal projectors are listed in Table 1. For a binary $[n,k,d]$ LCD code $C$, it is easy to see that $C^{\prime}=\{(0,c)\mid c\in C\}$ is a binary LCD $[n+1,k,d]$ code. Using this property, we obtained three optimal binary LCD codes from some codes listed in Theorem 4.2. Consequently, we obtain the following theorem:
@@
Any orthogonal projector mentioned above has constant diagonal entries, and hence any binary LCD code in Theorem 4.2 is either an even code or its dual code. Note that the orthogonal projectors of LCD codes above (or their dual codes) are adjacent matrices of regular graphs by construction.
@@ 5 Concluding Remarks
In this paper, we showed that binary even LCD codes and certain graphs are in one-to-one correspondence in a certain sense. Furthermore, we provided lower bounds on minimum weights of codes obtained from strongly regular graphs. Consequently, a part of the conjecture by Haemers, Peeters, Rijckevorsel [Haemers1999-no] was solved. The conjecture that the Paley graph gives a code with the largest weight among all the strongly regular graphs is still open for $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q\equiv 1\pmod{8}$ and $q>41$. Our result can be applied to generate a graph that gives optimal binary code, as shown in Section 4. We hope that studying such graphs leads to a discovery of a further relation between codes and graphs.

### v2

@@ Abstract
Keywords: LCD code, strongly regular graph.
Mathematical subject classification (2020): 94B05, 05C50.
@@ 1 Introduction
The interplay between codes and designs have provided many useful results [Assmus1992-rb]. Concerning relations between codes and graphs, several results are known: The rank of an adjacency matrix $A$ over $\mathbb{F}_{p}$ is called the $p$-rank of $A$. Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of strongly regular graphs are characterized by their eigenvalues. Massey [massey_linear_1992] introduced linear complementary dual codes (LCD codes for short), and showed that a linear code has the orthogonal projector if and only if it is an LCD code. Since any orthogonal projector is symmetric matrix, adjacency matrices of some undirected graphs should coincide with orthogonal projectors. Based on this observation, Rodrigues and Keys studied a relation between LCD codes and adjacency matrices of graphs. They gave two conditions where adjacency matrices of graphs generate LCD codes over $\mathbb{F}_{q}$, where $q$ is a prime power [Key2018-ec, Proposition 2].
In this paper, we consider binary LCD codes in particular, and study a relation between binary LCD codes and adjacency matrices of simple graphs. It is known that the orthogonal projector $A$ of a binary LCD code is a symmetric matrix with $A^{2}=A$, where we regard $A$ as a matrix over $\mathbb{F}_{2}$. Hence any adjacency matrix satisfies the latter condition is an orthogonal projector of binary LCD code, as shown in [Key2018-ec, Brouwer1992-nc]. We show that such adjacency matrix is the orthogonal projector of a binary even LCD code, and two binary even LCD codes are equivalent if and only if the corresponding graphs are isomorphic. This establishes a one-to-one correspondence between binary even LCD codes and the above-mentioned class of graphs. Haemers, Peeters and Rijckevorsel [Haemers1999-no] studied binary codes from adjacency matrices of strongly regular graphs. For strongly regular graphs with certain parameters, they observed that non-isomorphic graphs give inequivalent codes. This paper gives an explanation to the above observation. Furthermore, we improve some lower bounds on the minimum weight of binary LCD codes by constructing optimal binary LCD $[n,k,d]$ codes with the following parameters:
@@
We note that orthogonal projectors of the above optimal codes are adjacency matrices of regular graphs. By the above example, we show that from optimal binary even LCD codes we can generate graphs whose adjacency matrices give codes with large minimum weight.
@@ 2 Preliminaries
Let $\mathbb{F}_{q}$ be the finite field of order $q$, where $q$ is a prime power. Let $\mathbb{F}_{q}^{n}$ be the vector space of all $n$-tuples over $\mathbb{F}_{q}$. A $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ is said to be an $[n,k]$ code over $\mathbb{F}_{q}$. Especially, codes over $\mathbb{F}_{2}$ are said to be binary codes. Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$. The parameters $n$ and $k$ are said to be the length and the dimension of $C$, respectively. A vector in $C$ is called a codeword. The weight of $x=(x_{1},x_{2},\dots,x_{n})\in\mathbb{F}_{q}^{n}$ is defined as $\operatorname{wt}(x)=|\{i\mid x_{i}\neq 0\}|$. The minimum weight $d(C)$ of $C$ is the minimum weight among all nonzero codewords in $C$. If the minimum weight of $C$ equals to $d$, then $C$ is said to be an $[n,k,d]$ code over $\mathbb{F}_{q}$. We say that a code $C$ is even if all codewords have even weights; otherwise, the code $C$ is odd. Two $[n,k]$ codes $C_{1}$ and $C_{2}$ over $\mathbb{F}_{q}$ are equivalent if there is a monomial matrix $M$ such that $C_{2}=\{cM\mid c\in C_{1}\}$. The equivalence of two codes $C_{1}$ and $C_{2}$ is denoted by $C_{1}\simeq C_{2}$. A generator matrix of a code $C$ is any matrix whose rows form a basis of $C$.
@@
Let $C$ be a code over $\mathbb{F}_{q}$ with length $n$ and let $T\subseteq\{1,2,\dots,n\}$. Deleting every coordinate $i\in T$ in every codeword of $C$ gives a code called the punctured code on $T$ and denoted by $C^{T}$. Define a subcode $C(T)=\{(c_{1},c_{2},\dots,c_{n})\in C\mid c_{i}=0\text{ for all }i\in T\}$. Puncturing $C(T)$ on $T$ gives a code called the shortened code on $T$ and denoted by $C_{T}$. If $|T|=1$, say $T=\{i\}$, then we will write $C^{\{i\}}$ and $C_{\{i\}}$ simply as $C^{i}$ and $C_{i}$, respectively. Bouyuklieva gave a characterization of both punctured codes and shortened codes of binary LCD codes:
@@
The rank of an adjacency matrix $A$ over $\mathbb{F}_{p}$ is called the $p$-rank of $A$. Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of some strongly regular graphs are characterized by their eigenvalues. Haemers, Peeters and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs, and obtained several computational results for strongly regular graphs with small vertices. Keys and Rodrigues [Key2018-ec] constructed many LCD codes over $\mathbb{F}_{q}$ from adjacency matrices of strongly regular graphs via orthogonal projectors. Especially, they introduced the concept of reflexive LCD codes and complementary LCD codes [Key2018-ec, Definition 2], and provided the condition where adjacency matrices of graphs give LCD codes over $\mathbb{F}_{q}$ [Key2018-ec, Proposition 2].
@@ 3 Main results
In this section, we give main results on a correspondence between binary even LCD codes and adjacency matrices of some simple graphs. Also, we provide a lower bound on the minimum weights of binary codes that come from strongly regular graphs.
@@ 3.1 Characterization
Let $e_{i}$ be the unit vector whose components are all zero except $i$-th component, and let $r_{i}$ be the $i$-th row of $A$. Then, $Ae_{i}=r_{i}\in C$ for all $1\leq i\leq n$. Hence the row span of $A$ is a subspace of $C$. On the other hand, for any $c\in C$ it holds that $Ac=c$. This tells us that any codeword $c$ can be expressed as a linear combination of the rows of $A$. Hence $C$ is a subspace of the row span of $A$. This completes the proof. ∎
@@
  • If two vertices $x$ and $y$ are adjacent, then the number of the common neighbors of $x$ and $y$ is odd.
  • If two vertices $x$ and $y$ are not adjacent, then the number of the common neighbors of $x$ and $y$ is even.
@@
Bouyuklieva studied punctured codes and shortened codes of binary LCD codes and obtained Theorem 2.2 (see Section 2). To prove the theorem, they essentially considered the orthogonal projection of the unit vector $e_{i}$ and made the following observation: If the $i$-th component of the orthogonal projection of $e_{i}$ is $1$, then $C_{i}$ is LCD and $C^{i}$ is not LCD; otherwise, $C^{i}$ is LCD and $C_{i}$ is not LCD. The $i$-th component of the orthogonal projection of $e_{i}$ is equal to the $i$-th diagonal entry of the orthogonal projector $A$. Hence it follows that $(A)_{i,i}=0$ if and only if $C^{i}$ is LCD, and $A_{i,i}=1$ if only if $C_{i}$ is LCD.
@@ 3.2 Bounds on minimum weights
In this subsection, we give a lower bound on the minimum weights of binary codes obtained as row spans of adjacency matrices of strongly regular graphs. These results will be applied in Section 4. Note that all the assertions in this subsection hold not only for binary LCD codes but also for any binary codes.
@@
However, this is absurd since the left hand side of the equation is even while the right hand side is odd. This completes the proof. ∎
@@ 4 Applications
4.1 Binary codes from the Paley graph
Brouwer and Eijl [Brouwer1992-nc] proved that $p$-ranks of adjacency matrices of strongly regular graphs are characterized by their eigenvalues. For strongly regular graphs $srg(41,20,9,10)$, any corresponding binary code has $2$-rank, namely, dimension $20$. Haemers, Peeters and Rijckevorsel [Haemers1999-no] considered binary codes obtained as row spans of strongly regular graphs. They observed that for $srg(25,12,5,6)$ and $srg(41,20,9,10)$, any pair of non-isomorphic graphs gave a pair of inequivalent codes, and among such codes the code from the Paley graph had the largest minimum weight. Based on this result, they conjectured that the latter statement holds in general (see [Haemers1999-no, Section 4.3]). The classification of $srg(25,12,5,6)$ has been completed [paulus1973conference]; there are exactly $15$ non-isomorphic $srg(25,12,5,6)$, and with the aid of computer it was shown in [Haemers1999-no, Section 6.1] that the Paley graph gives the code with largest minimum weights.
On the other hand, the classification of $srg(41,20,9,10)$ has not been completed yet. However, it is not difficult to show the Paley graph of order $41$ gives a binary even $[41,20]$ code whose minimum weight is largest among all $srg(41,20,9,10)$. One way to see this is to use the Griesmer bound, which tells us the minimum weight of binary LCD $[41,20]$ code is at most $11$. By Lemma 3.6, the minimum weight must be even, and hence the result follows. According to [grassl_bounds_nodate], the largest minimum weight of binary linear code is exactly $10$, so it is obvious that the code from the Paley graph of order $41$ attains the largest minimum weight. To best of our knowledge, the conjecture is still open for $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q>41$ and $q\equiv 1\pmod{8}$.
Remarkably, the first part of the above observation by Haemers, Peeters and Rijckevorsel holds in general for any strongly regular graph with parameters $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ for $q\equiv 1\pmod{8}$ since any adjacency matrix of such a graph satisfies the condition given in Theorem 3.7. We state this fact as the following corollary:
Corollary 4.1.
Any pair of non-isomorphic $srg(q,(q-1)/2,(q-5)/4,(q-1)/4)$ with $q\equiv 1\pmod{8}$ gives a pair of inequivalent binary even LCD codes, and vice versa.
@@ 4.1 Binary codes from the Paley graph
By Proposition 3.10, any code from $srg(41,20,9,10)$ must have the minimum weight larger than $d\geq 1+(k+1)/(\min\{\lambda,\mu\}+1)=3$. Applying Theorem 3.7, we can improve the lower bound by $1$, and hence the lower bound on the minimum weight is $4$. The above observation shows this lower bound is tight for $srg(41,20,9,10)$.
4.2 Optimal LCD codes from regular graphs
In this subsection, we improve some lower bounds on the minimum weight of binary LCD codes by constructing optimal binary LCD codes from their orthogonal projectors. For recent results on optimal binary LCD codes, we refer the reader to [bouyuklieva_optimal_2021]. Let $d(n,k)$ denote the largest minimum weight among all LCD $[n,k]$ code.
@@ 4.2 Optimal LCD codes from regular graphs
For each of these codes, the first rows of their orthogonal projectors are listed in Table 1. For a binary LCD $[n,k,d]$ code $C$, it is easy to see that $C^{\prime}=\{(0,c)\mid c\in C\}$ is a binary LCD $[n+1,k,d]$ code. Using this property, we obtained three optimal binary LCD codes from some codes listed in Theorem 4.2. Consequently, we obtain the following theorem:
@@
Any orthogonal projector mentioned above has constant diagonal entries, and hence any binary LCD code in Theorem 4.2 is either an even code or its dual code. Note that the orthogonal projectors of LCD codes above (or their dual codes) are adjacent matrices of regular graphs by construction. It is remarkable that adjacency matrices of regular graphs can give LCD codes with the largest minimum weights. Furthermore, we remark that any code in Theorem 4.2 has odd length. This follows by the fact that the dimension of any binary even LCD code is even, and that the dual code of any even code must contain the all-ones vector; if a binary even LCD $[n,k]$ code $C$ has even length, then the all-ones vector must be in $C\cap C^{\perp}$, a contradiction.
@@ 5 Concluding remarks
In this paper, we proved that binary even LCD codes and some graphs are in one-to-one correspondence in a certain way. Furthermore, we provided lower bounds on minimum weights of codes obtained from strongly regular graphs. Consequently, we gave an explanation to the observation by Haemers, Peeters, Rijckevorsel [Haemers1999-no]. Our result can be applied to generate a graph that gives optimal binary code, as shown in Section 4. We hope that studying such graphs leads to a discovery of a further relation between graphs and even LCD codes.
6 Acknowledgement
The author would like to thank Stefka Bouyuklieva for several questions and comments that improved the manuscript.